A man is 20 yrs older than his son.In 10 yrs, he will be twice as old as his will be then be.How old is each?

Answer by checkley77(12844)

Son=x
Dad=x+20
(x+20)+10=2(x+10)
x+30=2x=20
x-2x=20-30
-x=-10
x=10 age of the son now.
10+20=30 age of the dad now.
Proof:
(10+20)+10=2(10+10)
30+10=2*20
40=40